# Reliable one-bit quantization of bandlimited graph data via single-shot noise shaping

### v1

@@ 1.1 Contribution
In this paper, we propose a novel single-shot noise shaping method for quantizing graph data. It draws inspiration from a recent result of Maly and Saab (2023) on neural network quantization. Similar to the approach of Krahmer et al. (2023, 2026), our method translates an $N$-dimensional graph structured signal into an $N$-dimensional quantized representation. In contrast to this existing method — which is computationally favorable, but restricted to using $\log\log(N)$ bits per entry — our approach allows flexible tuning of the bit-level, including the extreme case of using a single bit per data coefficient. Furthermore, it comes with slightly improved theoretical worst-case guarantees and a strongly simplified proof strategy.
@@ 1.2 Notation
In the following, we will abbreviate index sets as $N]=\{1,\dots,N\}$. We will use bold lowercase and uppercase letters to distinguish vectors and matrices from scalars. The Euclidean and the max-norm of a vector $\bm{z}\in\mathbb{R}^{N}$ are denoted by $\|\bm{z}\|_{2}$ and $\|\bm{z}\|_{\infty}$, respectively. The $\ell_{2}$-operator norm of a matrix $\bm{X}\in\mathbb{R}^{n_{1}\times n_{2}}$ is denoted by $\|\bm{X}\|$. The $(\ell_{2}\to\ell_{\infty})$-operator norm of $\bm{X}$ is denoted by $\|\bm{X}\|_{2,\infty}$. Note that $\|\bm{X}\|_{2,\infty}$ corresponds to the maximum $\ell_{2}$-norm of the rows of $\bm{X}$. When applied to a matrix $\bm{X}^{n\times n}$, the operator $\operatorname{diag}$ extracts its diagonal $\bm{x}\in\mathbb{R}^{n}$; when applied to a vector $\bm{x}\in\mathbb{R}^{n}$, $\operatorname{diag}$ returns the corresponding diagonal matrix $\bm{X}\in\mathbb{R}^{n\times n}$. We abbreviate $a\leq Cb$, for an absolute constant, by $a\lesssim b$ and write $a\simeq b$ if $a\lesssim b$ and $b\lesssim a$.
@@ 2 Preliminaries on graph data
Data on the graph can be represented by a function $f\!:\!\mathcal{V}\!\to\!\mathbb{R}$ defined on vertices of the graph. Note that any $f$ is equivalent to a vector $\bm{f}\in\mathbb{R}^{N}$}, where the $n$th component of $\bm{f}$ represents the function value at the $n$-th vertex in $\mathcal{V}$. The eigenvectors and eigenvalues of the Laplacian $\mathcal{L}$ provide a spectral interpretation of the graph. Indeed, the eigenvalues $\{\lambda_{1},\lambda_{2},\cdots,\lambda_{N}\}$ can be interpreted as graph frequencies, while the eigenvectors demonstrate increasing oscillatory behavior as the magnitude of the graph frequencies increases (Shuman et al., 2016). Based on this interpretation, the Graph Fourier Transform (GFT) of $\bm{f}$ is defined as $\widehat{\bm{f}}=\bm{X}^{T}\bm{f}\in\mathbb{R}^{N}$ with entries $\widehat{\bm{f}}(\lambda_{i})=\langle\bm{f},\bm{x}_{i}\rangle$.
In applications, graph data often exhibit intrinsic structure that can be leveraged for efficient data processing, e.g., dominance of low frequencies in the spectral domain (Dong et al., 2016). Graph data $\bm{f}$ are called bandlimited if there exists $r\in[N]$ such that the support of the GFT of $\bm{f}$ is contained in the frequencies $\{\lambda_{1},\dots,\lambda_{r}\}$ (Pesenson, 2009; Shuman et al., 2016).
Let us denote by $\bm{X}_{r}=[\bm{x}_{1},\cdots,\bm{x}_{r}]\in\mathbb{R}^{N\times r}$ the matrix $\bm{X}$ restricted to the first $r$ columns. Then every $r$-bandlimited function $\bm{f}$ can be written as $\bm{f}=\bm{X}_{r}{\bm{\alpha}}$, for some $\bm{\alpha}\in\mathbb{R}^{r}$}. By re-normalizing our data, we can always assume that $\|\bm{f}\|_{\infty}=1$}. The graph’s geometry and properties of its bandlimited function subspaces can be quantified by the graph incoherence (Shuman et al., 2016).
@@
(Shuman et al., 2016) For the $r$-dimensional subspace of graph functions spanned by $\bm{X}_{r}=[\bm{x}_{1},\cdots,\bm{x}_{r}]\in\mathbb{R}^{N\times r}$}, let $P_{\bm{X}_{r}}=\bm{X}_{r}\bm{X}_{r}^{T}$} be the orthogonal projection onto $\bm{X}_{r}$. The incoherence of the graph subspace $\bm{X}_{r}$ is defined as
$\mu(\bm{X}_{r})=\frac{N}{r}\max_{1\leq i\leq N}\|P_{\bm{X}_{r}}\bm{e}_{i}\|_{2}^{2}.$ (2.1)
It has been shown for various types of random graphs that $\mu$ is small with high probability if the respective graph is sufficiently large (Dekel et al., 2011; Brooks and Lindenstrauss, 2013). This indicates that the graph Laplacian eigenvectors are well-spread in general.
@@ 3.1 Single-shot noise-shaping (SSNS)
To quantize band-limited graph data $\bm{f}$, we will draw inspiration from recent work on neural network quantization (Maly and Saab, 2023). The authors propose to combine a simple neuron preprocessing algorithm with plain application of MSQ to quantize neurons in feedforward networks while preserving the input-output relation. It turns out that a similar strategy can be used for reliable quantization of bandlimited graph data. Compared to the iterative noise-shaping quantizer of Krahmer et al. (2023, 2026) this approach can be interpreted as a single-shot noise-shaping quantizer, which is also reflected in the name of our method.
Algorithm 1 Vector preprocessing (Maly and Saab, 2023)
@@
We recall the preprocessing algorithm of Maly and Saab (2023) tailored to the graph setting in Algorithm 1. The main goal of this preprocessing step is to obtain an alternative representation of $\bm{f}$ that preserves its low-frequency information while simultaneously minimizing the quantization error when followed by MSQ. Note that in Algorithm 1 we denote the kernel of $\bm{A}\in\mathbb{R}^{n_{1}\times n_{2}}$ restricted to $J\subset[n_{2}]$ by
@@
and that $\ker_{J_{k}}(\bm{X}_{k})\neq\{\bm{0}\}$ as long as $\||\bm{z}_{k}|-c\bm{1}\|_{0}>r$.
@@
In plain words: After preprocessing, all but at most $r$ entries of $\widehat{\bm{f}}$ lie exactly on the boundary of the unit $\ell_{\infty}$-ball and therefore attain the extreme values $\pm 1$. This property is crucial: when applying a midrise quantizer with an alphabet including $\pm 1$, at most $r$ unsaturated entries can contribute nonzero quantization error, while the remaining entries are already elements of the quantization alphabet. Figure 1 illustrates in $\mathbb{R}^{3}$ how Algorithm 1 proceeds.
@@
$\Gamma(\bm{X})=\sup_{\begin{subarray}{c}T\subset[N_{0}]\\ |T|=m\end{subarray}}\|\bm{X}^{T}|_{T}\|$
@@
$\leq\mu(\bm{X}_{r})\sqrt{\frac{r}{N}}\cdot\|\bm{\alpha}\|_{2},$
for some $\bm{\alpha}\in\mathbb{R}^{r}$ with $\|\bm{\alpha}\|_{2}=\|\bm{f}\|_{2}$, where we used that $\mu(\bm{X}_{r})\geq\frac{N}{r}\|\bm{X}_{r}\|_{2,\infty}$ by definition. Since this implies that $\|\bm{f}\|_{2}\geq\mu(\bm{X}_{r})^{-1}\sqrt{\frac{N}{r}}$, we obtain our claimed bound
@@ 3.2 Comparison with iterative noise-shaping methods
$\frac{\mathcal{Q}\mathcal{E}_{\bm{L}_{r}}(\bm{f},\bm{q})}{\|\bm{f}\|_{2}}\lesssim\mu(\bm{X}_{r})\frac{r\log(r)}{\sqrt{N\log(N)}}.$ (3.5)
It is important to note that this result requires a total bit budget of $N\log(\log(N))$ bits (Krahmer et al., 2023, 2026).
@@
Apparently, the theoretical error bound of our method surpasses the one of Krahmer et al. (2023, 2026). This is particularly remarkable as our approach is based on single-shot noise shaping before quantization while the work of Krahmer et al. (2023, 2026) adaptively shapes the noise during quantization. What is more, the comparison shows that the method of Krahmer et al. (2023, 2026) does not allow a final quantization level of only one bit per entry as soon as $\mathcal{A}$ is augmented in the final representation.
Our experiments on a ring graph in Section 4 suggest that the worst-case bound in Theorem 3.1 is tight. At the same time, they show that such worst-case guarantees might be too conservative in general since the empirical performance of the investigated methods considerably surpasses the predicted bounds on several graphs.
@@ 4 Numerical Experiments
While Algorithm 1 is well suited to convey the concept of our noise shaping approach, it requires $\mathcal{O}(r^{3}N)$ operations in the worst case. Meyer (2024) developed a more efficient implementation of Algorithm 1 reducing this to $\mathcal{O}(r^{2}N)$ operations. In our experiments, we will exclusively use the superior version of Meyer (2024), which is summarized in Algorithm 3 in the supplementary material, and refer to it as Single Shot Noise Shaping (SSNS).
(a) Grid graph
(b) Ring graph
(c) Swiss roll
(d) Bunny graph
Figure 4: Comparison of the performance of SSNS to the theoretical worst-case guarantees 3.1 and to the Step-by-Step-Serving with Replacement (SSS-R) method.
@@
We perform experiments with multiple bit budgets, $B\in\{1,2,4\}$, using $20$ independent realizations of the data for each bandwidth. SSNS is applied to each realization, and the quantization error is measured by the Relative Error in $\ell_{2}$-norm after applying a brick-wall filter or bandwith $r$, i.e.,
@@
Figure 2 shows the Relative Error of SSNS as a function of the bandwidth for each graph, on a semi-logarithmic scale and averaged over the random realizations. Each curve corresponds to a different bit budget ($B=1,2,4$). We observe the following.
  • As the number of bits increases, the average relative error decreases consistently across all graphs, confirming the expected trade-off between bit budget and quantization error. We further investigate this in Experiment 2.
@@
Experiment 2: Empirical and Theoretical Bit-Depth Scaling. In this experiment, we investigate the effect of quantization bit-depth on the reconstruction accuracy of bandlimited graph signals under the SSNS method, and compare its empirical performance with our theoretical error bounds. We consider the same graph topologies as in Experiment 1. For each graph, we generate bandlimited graph signals supported on the first $r=200$ graph Fourier modes, and generate $50$ independent signal realizations. We measure the relative squared reconstruction error (4.1) and report the average error over all realizations for each bit-depth.
Experiment 3: SSNS vs. SSS-R and Worst-Case Error Guarantees. Finally, we compare the performance of SSNS to our theoretical worst-case guarantees in Theorem 3.1 and to the Step-by-Step-Serving with Replacement (SSS-R) method of Krahmer et al. (2023, 2026) under a fixed quantization budget of $\log\log N$ bits required by the SSS-R method, cf. Section 3.2. Figure 4 reports the Relative Error (on logarithmic scale) as a function of $r$ for several representative graph topologies, including grid, ring, and Swiss-roll graphs. For each bandwidth value, the reported error is obtained by averaging over multiple random realizations.
@@
The dashed curves correspond to the theoretical upper bound in (3.6). In particular, we see that our theoretical bound is tight for the ring graph. For the remaining graphs, it is overly pessimistic due to its worst-case nature, but accurately captures the overall growth trend with respect to $r$ and provides theoretical justification for the observed degradation in performance at larger bandwidths. Importantly, SSNS remains well below the theoretical bound across all considered graphs, confirming its robustness to quantization effects.
(a) Original bunny shape
(c) Bunny shape, 1 bit SDW, $r=20$
(d) Bunny shape, 1 bit SSNS, $r=20$
(b) Bunny shape, 1 bit MSQ
(e) Bunny shape, 1 bit SDW, $r=50$
(f) Bunny shape, 1 bit SSNS $r=50$
Figure 6: Halftoning of the Stanford Bunny mesh graph signal using 1-bit quantization. (a) Original signal. (b) Memoryless scalar quantization (MSQ). (c),(e) Sigma-Delta-Weight (SDW) (Krahmer et al., 2023) with different bandwidth $r$. (d),(f) Proposed Single-Shot Noise Shaping (SSNS) with different different bandwidth $r$. All halftoned outputs are binary signals $\bm{q}\in\{-1,1\}^{N}$ visualized on the mesh.
Experiment 4: Graph-Signal Halftoning on 3D Meshes. Finally, we evaluate the proposed SSNS algorithm for halftoning of graph signals on 3D shapes and compare it with Sigma-Delta-Weight (SDW) in (Krahmer et al., 2023) and memoryless scalar quantization (MSQ); see also related work on 3D halftoning in (Gooran and Abedini, 2020, 2022; Mao et al., 2017; Lou and Stucki, 1998; Abedini et al., 2023; Abedini, 2023).
This experimental design follows the classical motivation of digital halftoning: although the displayed signal is binary, perceptual integration by the human visual system acts as an implicit low-pass filter, so visual appearance is primarily determined by low-frequency content (Floyd and Steinberg, 1976; Lyu and Wang, 2023; Krahmer and Veselovska, 2022). While the signal is not bandlimited and therefore lies outside the scope of our theoretical analysis, the three methods produce visually distinct halftoning patterns, illustrating different ways of shaping quantization artifacts on graphs. We emphasize that with higher bandwidth $r$, among the compared approaches, only SSNS is supported by provable error bounds, highlighting its role as a theoretically justified alternative for graph-based halftoning.
Experiments — Synopsis. The experiments confirm the effectiveness of our method for quantizing graph data with low bandwidth. The reduction in relative error with increasing bit budget is consistent across all tested graphs, demonstrating the method’s robustness to both signal structure and graph topology.
Furthermore, our results provide insights into the performance scaling of SSNS. Namely, graphs with more uniform connectivity (e.g., 2D grid) tend to exhibit smoother error growth compared to irregular graphs (e.g., sensor or bunny), reflecting the influence of graph topology on the distribution of Fourier energy.

### v2

@@ 1.1 Contribution
In this paper, we propose a novel single-shot noise shaping method for quantizing graph data. It draws inspiration from a recent result of Maly and Saab (2023) on neural network quantization. A key contribution of the present work is to show that the linear quantization framework of Maly and Saab (2023) can be lifted to graph domains by using incoherence to account for the underlying graph geometry. This lifting leads to the removal of restrictions on the realizable bit-levels of the quantization scheme, which existed in prior works on graph data quantization (Krahmer et al., 2023, 2026).
Our method translates an $N$-dimensional graph structured signal into an $N$-dimensional quantized representation. In contrast to the method of (Krahmer et al., 2023) — which is computationally favorable, but restricted to using $\log\log(N)$ bits per entry — our approach allows flexible tuning of the bit-level, including the extreme case of using a single bit per data coefficient. Furthermore, it comes with notably improved theoretical worst-case guarantees and a strongly simplified proof strategy.
@@ 1.2 Notation
In the following, we will abbreviate index sets as $N]=\{1,\dots,N\}$. We will use bold lowercase and uppercase letters to distinguish vectors and matrices from scalars. The Euclidean and the max-norm of a vector $\bm{z}\in\mathbb{R}^{N}$ are denoted by $\|\bm{z}\|_{2}$ and $\|\bm{z}\|_{\infty}$, respectively. The $\ell_{2}$-operator norm of a matrix $\bm{X}\in\mathbb{R}^{n_{1}\times n_{2}}$ is denoted by $\|\bm{X}\|$. The $(\ell_{2}\to\ell_{\infty})$-operator norm of $\bm{X}$ is denoted by $\|\bm{X}\|_{2,\infty}$. Note that $\|\bm{X}\|_{2,\infty}$ corresponds to the maximum $\ell_{2}$-norm of the rows of $\bm{X}$. We denote the kernel of $\bm{X}$ by $\ker(\bm{X})$. When applied to a matrix $\bm{X}^{n\times n}$, the operator $\operatorname{diag}$ extracts its diagonal $\bm{x}\in\mathbb{R}^{n}$; when applied to a vector $\bm{x}\in\mathbb{R}^{n}$, $\operatorname{diag}$ returns the corresponding diagonal matrix $\bm{X}\in\mathbb{R}^{n\times n}$. We abbreviate $a\leq Cb$, for an absolute constant, by $a\lesssim b$ and write $a\simeq b$ if $a\lesssim b$ and $b\lesssim a$.
@@ 2 Preliminaries on graph data
Data on the graph can be represented by a function $f\!:\!\mathcal{V}\!\to\!\mathbb{R}$ defined on vertices of the graph. Note that any $f$ is equivalent to a vector $\bm{f}\in\mathbb{R}^{N}$}, where the $n$th component of $\bm{f}$ represents the function value at the $n$-th vertex in $\mathcal{V}$. The eigenvectors and eigenvalues of the Laplacian $\mathcal{L}$ provide a spectral interpretation of the graph. Indeed, the eigenvalues $\{\lambda_{1},\lambda_{2},\cdots,\lambda_{N}\}$ can be interpreted as graph frequencies, while the eigenvectors demonstrate increasing oscillatory behavior as the magnitude of the graph frequencies increases Shuman et al. (2016). Based on this interpretation, the Graph Fourier Transform (GFT) of $\bm{f}$ is defined as $\widehat{\bm{f}}=\bm{X}^{T}\bm{f}\in\mathbb{R}^{N}$ with entries $\widehat{\bm{f}}(\lambda_{i})=\langle\bm{f},\bm{x}_{i}\rangle$.
In applications, graph data often exhibit intrinsic structure that can be leveraged for efficient data processing, e.g., dominance of low frequencies in the spectral domain Dong et al. (2016). Graph data $\bm{f}$ are called bandlimited if there exists $r\in[N]$ such that the support of the GFT of $\bm{f}$ is contained in the frequencies $\{\lambda_{1},\dots,\lambda_{r}\}$ Pesenson (2009); Shuman et al. (2016).
Let us denote by $\bm{X}_{r}=[\bm{x}_{1},\cdots,\bm{x}_{r}]\in\mathbb{R}^{N\times r}$ the matrix $\bm{X}$ restricted to the first $r$ columns. Then every $r$-bandlimited function $\bm{f}$ can be written as $\bm{f}=\bm{X}_{r}{\bm{\alpha}}$, for some $\bm{\alpha}\in\mathbb{R}^{r}$}. By re-normalizing our data, we can always assume that $\|\bm{f}\|_{\infty}=1$}. The graph’s geometry and properties of its bandlimited function subspaces can be quantified by the graph incoherence Shuman et al. (2016). [^1] Krahmer et al. (2023) worked with a squared incoherence, i.e., their incoherence parameter $\mu$ is equal to $\mu(\bm{X}_{r})^{2}$ as defined in (2.1). Note that our definition aligns with the original incoherence of Shuman et al. (2016) and leads to tighter bounds since $\mu(\bm{X}_{r})\leq\mu(\bm{X}_{r})^{2}\in[1,N/r]$, cf. Section 3.2 below.
@@
Shuman et al. (2016) For the $r$-dimensional subspace of graph functions spanned by $\bm{X}_{r}=[\bm{x}_{1},\cdots,\bm{x}_{r}]\in\mathbb{R}^{N\times r}$}, let $P_{\bm{X}_{r}}=\bm{X}_{r}\bm{X}_{r}^{T}$} be the orthogonal projection onto $\bm{X}_{r}$. The incoherence $\mu(\bm{X}_{r})$ of the graph subspace $\bm{X}_{r}$ is defined via
$\mu(\bm{X}_{r})^{2}=\frac{N}{r}\max_{1\leq i\leq N}\|P_{\bm{X}_{r}}\bm{e}_{i}\|_{2}^{2}.$ (2.1)
It has been shown for various types of random graphs that $\mu$ is small with high probability if the respective graph is sufficiently large Dekel et al. (2011); Brooks and Lindenstrauss (2013). This indicates that the graph Laplacian eigenvectors are well-spread in general.
@@ 3.1 Single-shot noise-shaping (SSNS)
To quantize band-limited graph data $\bm{f}$, we will draw inspiration from recent work on neural network quantization Maly and Saab (2023). The authors propose to combine a simple neuron preprocessing algorithm with plain application of MSQ to quantize neurons in feedforward networks while preserving the input-output relation. It turns out that a similar strategy can be used for reliable quantization of bandlimited graph data. Compared to the iterative noise-shaping quantizer of Krahmer et al. (2023, 2026) this approach can be interpreted as a single-shot noise-shaping quantizer, which is also reflected in the name of our method.
Algorithm 1 Vector preprocessing Maly and Saab (2023)
@@
We recall the preprocessing algorithm of Maly and Saab (2023) tailored to the graph setting in Algorithm 1. The main goal of this preprocessing step is to obtain an alternative representation of $\bm{f}$ that preserves its low-frequency information while simultaneously minimizing the quantization error when followed by MSQ. While not explicitly mentioned by Maly and Saab (2023), Algorithm 1 is closely related to the Gram-Schmidt walk of Bansal et al. (2018). Similar methods appeared far earlier in articles on “integer-making” theorems (Beck and Fiala, 1981) and balanced sampling (Deville and Tillé, 2004).
Figure 1 illustrates in $\mathbb{R}^{3}$ how Algorithm 1 proceeds. Intuitively, in each iteration $k$ it constructs a non-zero kernel element $\bm{b}_{k}$ of $\bm{X}$ that is zero in all entries in which the current iterate $\bm{z}_{k}$ already achieves the target $\ell_{\infty}$-norm $c$, and then adds a scaled version of $\bm{b}_{k}$ to set at least one entry of $\bm{z}_{k}$ to $\pm c$ while preserving the overall $\ell_{\infty}$-norm. Since in each step at least one entry of $\bm{z}_{0}$ is set to $\pm c$, Algorithm 1 terminates after at most NrN-r steps. Note that in Algorithm 1 we denote the kernel of $\bm{A}\in\mathbb{R}^{n_{1}\times n_{2}}$ restricted to $J\subset[n_{2}]$ by
@@
and that $\ker_{J_{k}}(\bm{X}_{k})\neq\{\bm{0}\}$ as long as $\||\bm{z}_{k}|-c\bm{1}\|_{0}>r$. Since $\ker_{J}(\bm{A})\subset\ker(\bm{A})$, all iterates $\bm{z}_{k}$ preserve the low-frequency information $\bm{X}_{r}\bm{z}_{k}=\bm{X}_{r}\bm{z}_{0}$ when Algorithm 1 is applied to $\bm{X}=\bm{X}_{r}$.
@@
After preprocessing, all but at most $r$ entries of $\widehat{\bm{f}}$ lie exactly on the boundary of the unit $\ell_{\infty}$-ball and therefore attain the extreme values $\pm 1$. This property is crucial: when applying a midrise quantizer with an alphabet including $\pm 1$, at most $r$ unsaturated entries can contribute nonzero quantization error, while the remaining entries are already elements of the quantization alphabet.
@@
Remark 3.2.
Theorem 3.1 is robust against spectral leakage in the following sense: If $\bm{f}$ is only of effective low bandwidth, i.e., $\bm{f}=\bm{f}_{r}+\bm{f}_{\text{high}}$ where $\bm{f}_{r}$ and $\bm{f}_{\text{high}}$ are the projections of $\bm{f}$ onto the $r$ first and all remaining frequencies, respectively, and we assume that $\|\bm{f}_{\text{high}}\|_{2}=\varepsilon\|\bm{f}\|_{2}$, for $\varepsilon\in(0,1)$, then Algorithm 1 could be applied to $\bm{f}_{r}$ and the bound in Theorem 3.1 would generalize to
$\frac{\mathcal{QE}_{L}(\bm{f},\bm{q})}{\|\bm{f}\|_{2}}\leq C\cdot 2^{-B}\cdot\mu(\bm{X}_{r})\frac{r}{\sqrt{N}}+\varepsilon.$
Note that our empirical results suggest robustness of the method to spectral leakage even when the algorithm is directly applied to $\bm{f}$ instead of $\bm{f}_{r}$.
@@
$\Gamma(\bm{X})=\sup_{\begin{subarray}{c}I\subset[N_{0}]\\ |I|=m\end{subarray}}\|\bm{X}^{T}|_{I}\|$
@@
$=\mu(\bm{X}_{r})\sqrt{\frac{r}{N}}\cdot\|\bm{\alpha}\|_{2},$
for some $\bm{\alpha}\in\mathbb{R}^{r}$ with $\|\bm{\alpha}\|_{2}=\|\bm{f}\|_{2}$, where we used that $\mu(\bm{X}_{r})^{2}=\frac{N}{r}\|\bm{X}_{r}\|_{2,\infty}^{2}$ by definition. Since this implies that $\|\bm{f}\|_{2}\geq\mu(\bm{X}_{r})^{-1}\sqrt{\frac{N}{r}}$, we obtain our claimed bound
@@ 3.2 Comparison with iterative noise-shaping methods
$\frac{\mathcal{Q}\mathcal{E}_{\bm{L}_{r}}(\bm{f},\bm{q})}{\|\bm{f}\|_{2}}\lesssim\mu(\bm{X}_{r})^{2}\frac{r\log(r)}{\sqrt{N\log(N)}}.$ (3.5)
It is important to keep in mind that this result requires a total bit budget of $N\log(\log(N))$ bits and that Krahmer et al. (2023, 2026) work with the squared incoherence $\mu(\bm{X}_{r})^{2}$, cf. Footnote 1.
@@
Apparently, the theoretical error bound of our method surpasses the one of Krahmer et al. (2023, 2026) in all relevant dependencies ($\mu(\bm{X}_{r})$, $r$, and $N$). This is particularly remarkable as our approach is based on single-shot noise shaping before quantization while the work of Krahmer et al. (2023, 2026) adaptively shapes the noise during quantization. What is more, the comparison shows that the method of Krahmer et al. (2023, 2026) does not allow a final quantization level of only one bit per entry as soon as $\mathcal{A}$ is augmented in the final representation.
Our experiments in Section 4 suggest that the worst-case bound in Theorem 3.1 might still be too conservative in general since the empirical performance of the investigated methods considerably surpasses the predicted bounds on several graphs.
@@ 4 Numerical Experiments
While Algorithm 1 is well suited to convey the concept of our noise shaping approach, it requires $\mathcal{O}(r^{3}N)$ operations in the worst case. Meyer (2024) developed a more efficient implementation of Algorithm 1 reducing this to $\mathcal{O}(r^{2}N)$ operations. In our experiments, we will exclusively use the superior version of Meyer (2024), which is summarized in Algorithm 3 in Appendix A.1, and refer to it as Single Shot Noise Shaping (SSNS).
Our experiments were conducted on a MacBook Pro equipped with an Apple M1 processor and 16 GB RAM, using MATLAB R2023a. Graph signal processing operations, including graph Fourier transform computations and graph visualizations, were implemented using the GSPBOX toolbox Perraudin et al. (2014). The corresponding code is available in the supplementary materials.
Figure 4: Comparison of the performance of SSNS to the theoretical worst-case guarantees 3.1, Memory-less quantization (MSQ), and to the Step-by-Step-Serving with Replacement (SSS-R) method.
@@
We perform experiments with multiple bit budgets, $B\in\{1,2,4\}$, using $20$ independent realizations of the data for each bandwidth. SSNS is applied to each realization, and the quantization error is measured by the Relative Error in $\ell_{2}$-norm after applying a brick-wall filter or bandwidth $r$, i.e.,
@@
Figure 2 shows the Relative Error of SSNS as a function of the bandwidth for each graph, on a semi-logarithmic scale and averaged over the random realizations. Each curve corresponds to a different bit budget ($B=1,2,4$). We observe the following:
  • As the number of bits increases, the average relative error decreases consistently across all graphs, confirming the expected trade-off between bit budget and quantization error. We further investigate this in Experiment 3 in Appendix A.2.
@@
Experiment 2: SSNS vs. SSS-R and Worst-Case Error Guarantees. In the second experiment, we compare the performance of SSNS to our theoretical worst-case guarantees in Theorem 3.1, to naive MSQ directly applied to the data, and to the Step-by-Step-Serving with Replacement (SSS-R) method of Krahmer et al. (2023, 2026) under a fixed quantization budget of $\log\log N$ bits required by the SSS-R method, cf. Section 3.2.333We did not include LGFF-based quantization (Reingruber and Matz, 2025) in Experiment 2. Their method is hard to compare directly to ours since it produces overcomplete representations $\bm{q}\in\mathcal{A}^{M}$, for $M\geq N$, and operates in the transform domain with bit allocation, where informative coefficients receive more bits, while less relevant coefficients receive fewer bits. Figure 4 reports the Relative Error (on logarithmic scale) as a function of $r$ for several representative graph topologies, including grid, ring, and Swiss-roll graphs. For each bandwidth value, the reported error is obtained by averaging over multiple random realizations (95% confidence intervals are provided).
@@
The dashed curves correspond to the theoretical upper bound in (3.6). In particular, we see that our theoretical bound is overly pessimistic due to its worst-case nature, but captures the overall growth trend with respect to $r$ and provides theoretical justification for the observed performance degradation at larger bandwidths. Importantly, SSNS remains well below the theoretical bound across all considered graphs, confirming its robustness to quantization effects.
Experiments — Synopsis. All conducted experiments confirm the effectiveness of our method for quantizing graph data with low bandwidth. The reduction in relative error with increasing bit budget is consistent across all tested graphs, demonstrating the method’s robustness to both signal structure and graph topology. Furthermore, our results provide insights into the performance scaling of SSNS. Namely, graphs with more uniform connectivity (e.g., 2D grid) tend to exhibit smoother error growth compared to irregular graphs (e.g., sensor or bunny), reflecting the influence of graph topology on the distribution of Fourier energy. Additional experiments are provided in Appendix A.2, including an application of our method to 3D digital halftoning, and comparisons of the runtime and the empirical and theoretical bid-depth scaling of the error.